Calculate the mass (g) of bromocyclohexane produced given that 1.56 mL of cyclohexane (density 0.779g/mL) is reacted with 1.80 mL of 0.87 M bromine. Give your answer to 3 places after the decimal.


1
Expert's answer
2021-03-29T06:06:46-0400

1) m(cyclohexane)= 0.779*1.56=1.215 g

2) n(cyclohexane)=1.215/84=0,0145 mole

3)n(Br2)=1.8*0,001*0.87=0,001566 mole

4)m(bromocyclohexane)= 0,001566*(80+12*6+11)=0,255g

answer: 0,255g of bromocyclohexane